A water tank is h feet high. Water flows at velocity v feet per second. v=6(sqrt)h
a. Find the velocity of water flow of 60 feet per second
b. find the velocity of water flow when hieght is 130 feet.
Round answer to two decimal places.
1
Expert's answer
2011-09-18T16:11:34-0400
a. H= (v/6)2=(60/6)2=100 feets.
b. V= 6(130)1/2=6*11.40=68.41feet per second.